# MToP: A MATLAB Optimization Platform for Evolutionary Multitasking

### 13 Dec 2023 v1

@@ IV-B Adding New Problems
Similar to adding algorithms, adding problems in MToP requires the creation of a specific problem class file that inherits the base class Problem.m in the Problems/ folder. The suggested maximum function evaluations maxFE need to be predefined in the constructor of the specific problem. After that, refer to Listing LABEL:list:prob to implement the configuration of each optimization task in setTasks() function. When optimization tasks include multi-objective optimization problems, it is necessary to additionally implement the getOptimum() function to return the optimum solution set or reference points. Finally, for setting customized parameters that can be modified in the GUI, it is also required to implement getParameter() and setParameter() functions.
@@ IV-C Adding New Performance Metrics
Users can add new performance metrics by creating a new metrics function file in the Metrics/ folder. Follow the pattern as in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object. The metric results displayed in the GUI table need to return the RowName, ColumnName, and TableData fields. The RowName for combined multitask metrics is typically the problem name, while metrics on individual tasks need to return the name of each task. The ColumnName is typically the algorithm name. The calculation function of TableData needs to be implemented for the specific performance metric. To implement the function to plot the convergence of the metric, the ConvergeData field is also required. Additionally, for multi-objective optimization, the metric function should also return non-dominated solutions in the ParetoData field.
@@ V Conclusion and Future Work
This paper introduced an open-source MATLAB platform MToP for EMT. MToP provides a user-friendly GUI, enriched algorithms and problems, and convenient code patterns. The current version of MToP includes more than 30 MTEAs, more than 30 single-task EAs (that can handle MTO problems), more than 150 MTO benchmark problems, and several real-world applications of EMT. The functional modules, file structure and code pattern of MToP are described in detail. Moreover, the procedures and extensions for conducting experiments using MToP are also presented.
The current MToP, while carefully reimplementing, modifying, and testing the included algorithms and problems, might still have some reproducing errors and bugs. We have been actively looking for and fixing the issues and continue to update MToP. Fortunately, MToP is uploaded on the open-source website GitHub, where followers and researchers can test and give feedback on issues. After the upload, we also received and updated many codes from researchers.
In the future, we will continue to develop and update MToP with inspiration and contributions from the academic community. We hope that MToP can become a convenient tool to help researchers advance the field of EMT and consequently advance the evolutionary computation field.

### 09 Apr 2024 v2

@@ IV-B Adding New Problems
To integrate new problems into MToP, users should create a dedicated problem class file within the Problems/ folder. This class file should inherit from the base class Problem.m. It’s advisable to define the suggested maximum function evaluations (maxFE) within the constructor of the specific problem. Subsequently, users can refer to Listing LABEL:list:prob for guidance on configuring each optimization task within the setTasks() function. In cases where optimization tasks involve multi-objective optimization problems, users need to additionally implement the getOptimum() function to provide the optimal solution set or reference points. Lastly, for setting customizable parameters that can be modified within the GUI, users must also implement the getParameter() and setParameter() functions.
@@ IV-C Adding New Performance Metrics
To incorporate new performance metrics into MToP, users should create a new metrics function file within the Metrics/ folder. Follow the structure outlined in Listing LABEL:list:metric to receive the data structure object MTOData and return the calculated result object. For the metrics displayed in the GUI table, ensure that the RowName, ColumnName, and TableData fields are properly populated. Typically, the RowName for combined multitask metrics is the problem name, while metrics on individual tasks should return the name of each task. The ColumnName typically represents the algorithm name. Implement the calculation function for the specific performance metric in the TableData field. For plotting metric convergence, ensure the ConvergeData field is populated. In the case of multi-objective optimization, include non-dominated solutions in the ParetoData field.
@@ V Conclusion and Future Work
This paper introduces MToP, an open-source MATLAB platform designed for EMT. MToP features a user-friendly GUI, a rich collection of algorithms and problems, and convenient code structures. The current version of MToP contains over 40 MTEAs, more than 40 single-task EAs capable of handling MTO problems, over 150 benchmark MTO problems, and several real-world applications of EMT. This paper provides detailed descriptions of MToP’s functional modules, file structure, and code patterns, as well as guidelines for conducting experiments using MToP.
While MToP has undergone careful reimplementation, modification, and testing of included algorithms and problems, it may still contain some reproducing errors and bugs. Continuous efforts are underway to identify and rectify such issues, with updates regularly posted on GitHub. The platform encourages feedback and contributions from followers and researchers, with many codes already received and integrated from the community.
Moving forward, the development and enhancement of MToP will continue, drawing inspiration and contributions from the academic community. It is our aspiration that MToP evolves into a valuable tool to propel research in the field of EMT and, by extension, advance the broader evolutionary computation field. Specifically, MToP is considering the incorporation of multitask combinatorial optimization problems [5, 6] and more real-world applications [8, 111] in the future.
@@ Acknowledgement
All algorithm, problem, and metric implementations within MToP adhere closely to the methodologies outlined in their respective articles. The code provided by the original authors undergoes modifications based on the source code during its integration into MToP.